Over a time interval of 1.88 years, the velocity of a planet orbiting a distant star reverses direction, changing from +19.0 km/s to -17.4 km/s. Find the following values, include the correct algebraic sign with your answers to convey the directions of the velocity and the acceleration. (a) the total change in the planet's velocity (b) its average acceleration during this interval
Answer #1

a) change in velocity = - 17.4 - 19 = - 36.4 km/s

b) a = change in velocity / time

a = - 36.4*1000/ (1.88* 365*24*3600)

a = - 6.14*10^-4 m/s^2

=========